OpenStudy (anonymous):

Tickets to the school play cost $2.50 for general admission of $2.00 for students. On opening night 319 tickets were sold. The total receipts were $694. How many of each kind of ticket were sold?

OpenStudy (campbell_st):

you have 2 equations that will need to be solved simultaneously let g = general admission and s = students Receipts: 2.5g + 2s = 694 attendance: g + s = 319 the easiest way to solve is but substitution. rewrite the attendance equation with s as the subject and substitute it into the 1st equation. Then you will be able to solve for g...

OpenStudy (anonymous):

ok but how would I rewrite the equation?

OpenStudy (campbell_st):

ok... so the attendance will become s = 319 - g substituting 2.5g + 2(319 - g) = 694 you'll need to distribute and collect like terms before solving for g.

OpenStudy (anonymous):

ok thanks